I need a tutor to help me with problem please below?
The number N of aluminum cans used each year varies directly as the numbe of people P using the cans. If 58 people use 14,964 cans in one year, how many cans are used in a city which has a population of 1,709,000?The number of cans used is ___ ?

Answer by stanbon(75887)

The number N of aluminum cans used each year varies directly as the number of people P using the cans.
N = kP
---
Solve for "k" using "58 people use 14,964 cans in one year":
14964 = k*58
k = 258
----
Equation:
N = 258*P
---
How many cans are used in a city which has a population of 1,709,000?
N = 258*1709000
N = 440,922,000
=======================
Cheers,
Stan H.
=======================